1. Let Q(x) denote the statement “x is an integer”. What are the truth values?



A. Q(-1)



B. Q(0)



C. Q(8/2)



D. Q(sqrt(-4))



E. Q(sqrt(4))




1
Expert's answer
2022-04-01T04:38:53-0400

a. true. -1 is integer.

b. true. 0 is integer

c. true 8/2=4 is integer

d. false. "\sqrt{-4}" is not integer

e.true "\sqrt{4}=2" is integer